OpenStudy (anonymous):

If a line passes through the points (-2, -4) and (3, -2), the equation of the line is y + 2 = 2/5(x - _____).

OpenStudy (anonymous):

y + 2 = 2/5(x -(-4))= 2/5(x+4))